8-pack control system working group Meeting Notes

3 April 2003


No meeting next week (Ron and Janice gone)



-On going: Need to check out readback with real signals, check archiving of variables (pwr, pmin, pmax, etc), and check the plots. No change.



-nothing new



DAC/ADC Issues

-Fixed? Some of the phases are clamped to +-180 and others are wrapped around (e.g. 200 goes to –160).  Kukhee will fix these to all wrap around; no clamping. Done.


Some displays need a precision of 5 to get gains to show correctly. For Kukhee.


All of these issues are for later:


-TBD: The trickle down of the SLED variables.  This was specified at a previous meeting (numbers on SLED panel need to trickle down to I/Q and A/P PVs).  For processing, we might want this feature functional sooner rather than later. 

-TBD: the trickle down of the variables on the power distribution panel as well.  There are 4 PVs there called Klystron Sled-in phases.  These need to trickle down to the phase DACs as well.  This phase is to be added to the channel phase angle (IQ calib panel) and to the timestep phase before going out to the DAC.



-JN report: Does 8-pack need a backup hardware trip circuit? 

Marc says yes.  NLCTA ops will take care of it.  Will need SW support for a fault counter, status and reset.  Marc thinks we’ll use Joe’s interlock at least for now.


-This is on hold for now:  SWE will make a PV for a temporary DIO channel as a reset to Joe Frisch’s kludge RE Interlock.  They will send the name to Janice who will put a button on the panel with the pulse width and the klystron attenuation.  Its DIO channel assignment should be on Andrew’s previously sent list.  Also need a channel for status when it trips off.  Judy to check for one in and one out available on DIO.

HW reset counter needed as well.

-(no change) The following item was deferred for a few weeks:

LLRF + NLCTA Ops need to sort out what the LLRF enable and the mod off buttons do


SCP Configs

-(no change) LLRF folks need to check mini-dbedit files for TRBRs and report problems if any



-All seem to work.  Folks should report problems as they arise.

-(no change) We deferred the following: longer term issue: what to do if reading out an SIS when a fault occurs?

-(no change) Right now the SISs are paused and read out with a user request only.  A read when a fault happens will also be added to this.



-Sooner than later Janice will check current list of archived variables to see if any are missing.  (Notably PAVE, PMAX, PMIN, PWR and the processing control).

-Status? The matlab archive retrieval functions seem to work.  Bob is working on compiling these so no matlab licenses are needed. Compiled version works now. 


Archiving of SISs

-(no change) SWE plans to implement the following two features requested by LLRF (Joe’s email):

have waveforms archived on (1) every trip and (2) when an acquire is done by a user

-(no change) Later SWE will work with Joe on his other two requests: archive every few minutes and when a change occurs in a selected set of values in EPICS (a pseudo-trip if you will)



-nothing new 



Still not found: Andrew says a panel exists from SIS data to check out these calibration numbers.  Janice will check. 

If not:(No change) All of the following will get going in the next few weeks:

-Each IQA channel has 6 or 7 calibration numbers that are module dependent.  Gain/offset I,Q, quadrature angle and overall phase.  These numbers are used to change the SIS waveforms from I/Q to A/P.  These numbers will go on the IQA module panel. 

-Steve will email the equations for this calculation (Mail arrived after the meeting)

-Kukhee needs to change the SIS converter subrecord and new PVs need to be created.

-Once names are decided upon, Janice can make the panel (Judy may have started this already.)

-These 6 (or 7) numbers x 8 channels x 4 IQAs need to be archived at least once every 10 minutes.  The panel interface could just be one table for each IQA with 8 rows for the 8 channels and 6 columns with the calibration variables.


Processing algorithm

- Processing algorithm ready to be checked.  Andy will call Kukhee and Ron when there’s a good time to check it. Processing needs to reset IQA too.


Peak Power Meter (PPM)

-Sliding: Plans for this week  (needed by Thursday 27th)

-Tonee and Janice will hook up enet box(es) and gpib cables to the PPMs.

-(no change) Tonee will tweak the labview program to send the waveform to epics once we have names – probably late next week

-(no change) Tonee and Janice will set up the necessary hardware in ESB (Enet box, cables, etc)



-(no change) We’re currently waiting for Dick’s setup to settle down then we’ll do one more iteration of getting list, changing database, changing panels, programming the dcm, changing alarm handler and archiver stuff as needed



-Are we still confused with out vs sled?  Yes.  Janice will look.

-Is this resolved?  Steve will try to use Ron’s workaround for getting epics to printout on the color printer near his office.  Either with the vax queue name (SLC:Qname) or the NT queue name (something similar).

-(no change)Judy reported the crate profiles on the web are updated and look good.  Janice requested that channel assignment information also be available via the web, even it it’s just a link to the template file.  Judy and Ron will look into this.

-Has this been done?  Andrew will tell Kukhee when he takes out the DIO module to modify it for the deadman

-(no change) we deferred the following:

-public/private matlab issues to be resolved at a later date
(Part of separate ongoing SW project)

-LLRF needs to decide what the reset button does.